# Pretraining Scaling Laws for Generative Evaluations of Language Models

### v1

@@ 4.2 Predicting Pass Rates from Model Parameters and Pretraining Tokens
Figure 6: Scaling of Pass Rates with Gold Reference Likelihood (Full Fit). Each panel fits Eqn. 6, $\log(\mathrm{pass}_{\mathcal{B}}@k)=\xi_{0}+K_{0}(k)\cdot[-\log\!\big(\mathrm{GoldProb}_{\mathcal{B}}\big)]^{\kappa(k)}$, for $k\in\{1,10^{2},10^{4}\}$. Regressing the gold reference log-likelihoods against the pass rates produces a visually tighter fit with less residual scatter than the compute scaling law (cf. Fig. 1). As $k$ increases, the irreducible error $\widehat{\xi}_{0}(k)$ falls toward zero and the exponent $\widehat{\kappa}(k)$ increases, making the relationship a steeper, purer power law.
@@ 5 Fitting and Predicting Pass Rates from Gold Reference Likelihoods
Generative evaluations oftentimes contain not just problems and correct answers, but also “gold references”: high quality responses that reach the correct answer from the given problem. For autoregressive language models, one can straightforwardly compute the likelihood of the gold reference given the problem, as well as the arithmetic mean over the benchmark’s gold reference probabilities:
$\mathrm{GoldProb_{\mathcal{B}}}\quad\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}\quad\frac{1}{|\mathcal{B}|}\,\sum_{i\in\mathcal{B}}\,p_{\theta}(\text{Gold Reference}_{i}|\text{Problem}_{i}),$ (5)
For our gold reference likelihood scaling law, we assessed how well the benchmark $\mathrm{pass_{\mathcal{B}}@k}$ can be (1) fit and (2) predicted as a function of the log likelihoods of the gold references.
@@ 5.1 Fitting Pass Rates from Gold Reference Likelihoods
Motivated by Schaeffer et al. (2024), we tested whether this average likelihood of the gold responses can be accurately regressed to fit and predict benchmark pass rates $\mathrm{pass_{\mathcal{B}}}@k$.
@@
Perhaps surprisingly, we discovered that this straightforwardly computable quantity serves as an accurate predictor of benchmark pass rates. As shown in Figure 6, regressing against the negative log likelihood of the gold references produces a visually tighter fit than the pretraining compute law (Fig. 1). The residual scatter is reduced, and the relationship between $\log(\mathrm{pass_{\mathcal{B}}}@k)$ and $\log(\mathrm{GoldProb_{\mathcal{B}}})$ has a smaller reducible error and more closely resembles a power law. We next examined how the number of attempts per problem $k$ influences the scaling law’s parameters (Fig. 11). Consistent with our findings for the prior two scaling laws, the irreducible error term $\xi_{0}(k)$ collapses toward zero as $k$ increases and the exponent $\kappa(k)$ increases smoothly.
Figure 7: Predicting Pass Rates from Gold Reference Likelihoods (Backtesting). While the overall predictive error is comparable to the other laws, it is slightly lower for small $k$ and slightly higher for large $k$. Top Left: The most striking result is the stability of the scaling law’s parameters. In stark contrast to the compute and parameters+tokens laws, whose parameters stabilized only when fit on models within $\mathord{\sim}2$ orders of magnitude of the target, the backtested estimates for $\widehat{\xi}_{0}(k)$, $\widehat{K}_{0}(k)$, and $\widehat{\kappa}(k)$ converge to their final values using models up to $\mathord{\sim}5$ orders of magnitude cheaper than the target. Other Three Panels: This remarkable stability suggests that the relationship between gold reference likelihoods and pass rates provides a robust signal, enabling reliable, long-range forecasts from much cheaper models.
@@ 5.2 Predicting Pass Rates from Gold Reference Likelihoods
Using our backtesting approach once more, we evaluated how well the gold references scaling law can predict the performance of our most expensive model. For small $k$, the gold reference likelihoods scaling law has slightly lower relative error at predicting the most expensive model’s performance, but worsens for larger $k$ (Fig. 7 Upper Left). The most remarkable result, however, is the stability of the scaling law’s parameters. In stark contrast to the compute and parameter + token scaling laws whose parameters only stabilized when fit on models within $\mathord{\sim}2$ orders of magnitude of the target, the parameters for the gold reference law are exceptionally robust: the backtested estimates for the three parameters converge to their final values when fitting on models up to $\mathord{\sim}5$ of magnitude cheaper than the target (Fig. 7, Other Three Panels) and for all values of $k$. This significantly stable finding suggests that the relationship between gold reference likelihoods and pass rates is a consistent signal, enabling long-range forecasts from comparatively inexpensive models.
6 How the Compute Scaling Law Emerges from the Parameters + Tokens Scaling Law
The approximately equivalent performance between the compute scaling law (Eqn. 3) and the parameters + tokens scaling law (Eqn. 4) led us to ask whether a deeper connection might exist. We discovered that the compute scaling law is the compute-optimal envelope of the parameters+tokens scaling law. We summarize the insight here and defer a full explanation to Appendix D.
Fix a benchmark $\mathcal{B}$ and attempts per problem $k$. Consider a compute budget $C\approx c\,ND$. The best achievable error is obtained by minimizing the right-hand side of Eqn. 4 over all $(N,D)$ with $ND=C/c$. Evaluating at the optimizer yields Eqn. 3. The mapping of exponents and constants is:
$\alpha(k)=\Bigl(\tfrac{1}{\beta(k)}+\tfrac{1}{\gamma(k)}\Bigr)^{-1},\qquad E_{0}(k)=\mathcal{E}_{0}(k),$
with a closed-form $C_{0}(k)$.
Compute-Optimal Allocation. Along the envelope (i.e., when Eqn. 3 is tight), the optimal split of compute between parameters and tokens is
@@ 6 How the Compute Scaling Law Emerges from the Parameters + Tokens Scaling Law
Thus the parameter/token ratio shift with scale according to $\beta,\gamma$; when $\beta=\gamma$, both grow like $C^{1/2}$.
Departures from the Envelope. If training does not follow the compute-optimal allocation $(N^{*},D^{*})$, define the misallocation ratio $r(C)=\tfrac{N}{N^{*}(C)}=\tfrac{D^{*}(C)}{D}$. The departure from the envelope is multiplicative on top of Eqn. 3:
$\log\big(\operatorname{pass}_{\mathcal{B}}\text{@}k\big)(C,k)=E_{0}(k)+C_{0}(k)\cdot C^{-\alpha(k)}\cdot\underbrace{\Bigl[\tfrac{\gamma}{\beta+\gamma}\,r^{-\beta}+\tfrac{\beta}{\beta+\gamma}\,r^{\gamma}\Bigr]}_{\displaystyle\Phi(r;\beta,\gamma)\ \geq 1},$
with equality iff $r\equiv 1$. Small deviations are benign ($\Phi=1+\tfrac{\beta\gamma}{2}\,(\log r)^{2}+O(\log r)^{3})$ ), but persistent off-ridge scaling yields a growing penalty. A useful case: if one scales $N$ and $D$ at a fixed ratio (i.e., without re-optimizing with $C$), the effective compute slope degrades to $\alpha_{\text{path}}=\frac{\min\{\beta,\gamma\}}{2}$.
Takeaways.
(i) The compute law is not an independent phenomenology; it is the optimal-allocation shadow of the parameters + tokens law. (ii) Use the allocation above to stay on-envelope when planning larger models. (iii) Implementation constants relating $C$ to $ND$ shift $C_{0}(k)$, but do not change $\alpha(k)$. All statements hold pointwise in $k$.
@@ 7 Discussion
In this work, we investigated pretraining scaling laws for generative evaluations of language models. Our work was motivated most directly by OpenAI et al. (2024)’s success at predicting pass rates at HumanEval (Chen et al., 2021), as well as by Hu et al. (2024).
Key Findings Our work’s main contribution is the proposal and rigorous backtesting of three distinct scaling laws for pass-at-$k$, using (1) pretraining compute, (2) model parameters and pretraining tokens, and (3) gold reference log likelihoods as covariates. We demonstrate that a key hyperparameter in generative evaluation – the number of attempts per problem $k$ – offers a novel lever for controlling the scaling law parameters and predictability. Empirically, we find that while all three laws have comparable predictive accuracy, the parameters of the gold reference likelihood law are uniquely stable, converging across nearly five orders of magnitude of compute. Finally, we establish a theoretical connection between the compute law and the parameters + tokens law, proving that the compute law emerges as the compute-optimal envelope of the parameters-and-tokens law.
Limitations The main limitation of this work is that the empirical results are based on a single benchmark (GSM8K) and a single model family (Pythia). Additionally, we consider only pass-at-$k$ generative evaluations, whereas other are of course worth considering as well.
Related Work Due to space constraints, we defer related work to Appendix A.
Future Directions Multiple next steps are clear. First, subsequent work should determine whether predictions can be made more accurately using either additional sampling or controlled pretrained models. Second, OpenAI et al. (2024) claimed subsetting and/or stratifying problems enabled better predictions; what exactly did this approach entail, and how much did it help? Third, the utility of the gold reference log likelihoods was perhaps surprising because if we think of next-token sampling as a branching process, there are likely exponentially many paths from the problem to the correct answer, and it is unclear that the benchmark-provided gold references have any meaningful relationship with likely samples from the model; this raises at least two future questions: (i) how general is this relationship across benchmarks and models, and (ii) to what extent can this indicator be trusted for predicting the performance of more expensive models, especially under optimization pressure?
@@ Appendix A Related Work
The study of neural scaling laws, which characterize how model performance predictably improves with resources, has become a cornerstone of modern large-scale machine learning (Hestness et al., 2017; Kaplan et al., 2020a; Hoffmann et al., 2022a; Roberts et al., 2022; Bahri et al., 2024). Foundational work demonstrated that pretraining loss follows a power law with respect to model size, dataset size, and training compute (Kaplan et al., 2020a; Henighan et al., 2020; Hoffmann et al., 2022a). These principles now guide the development of nearly all state-of-the-art language models (OpenAI et al., 2024; Team et al., 2024; Grattafiori et al., 2024; DeepSeek-AI et al., 2025; Qwen et al., 2025; Zhipu-AI et al., 2025). This research area has since expanded to explore scaling properties in diverse domains, including computer vision (Zhai et al., 2022), machine translation (Ghorbani et al., 2021), and even for optimizing inference-time compute rather than training (Wu et al., 2024; Snell et al., 2024). Despite this breadth, scaling laws for complex, multi-step generative tasks—which often represent the ultimate goal of language modeling—have remained comparatively underexplored due to the challenges of evaluation. Our work builds most directly on recent efforts to forecast performance on such generative evaluations. The GPT-4 Technical Report pioneered this direction by showing that the negative log pass rate on a subset of the HumanEval coding benchmark scaled predictably as a power law of pretraining compute (OpenAI et al., 2024). They successfully predicted GPT-4’s performance by extrapolating from smaller models over approximately three orders of magnitude (OpenAI et al., 2024). We adopt their proposed functional form for our compute-based law and extend their work by systematically analyzing how the number of attempts, $k$, reshapes the law’s parameters. Furthermore, our rigorous backtesting reveals the brittleness of this compute-only law, finding it reliable only across 1-2 orders of magnitude in our setting. Separately, our work shares a philosophical motivation with Hu et al. (2024), who investigate emergent abilities. They argue that seemingly unpredictable jumps in performance on discriminative tasks can be understood as smooth, continuous improvements when measured by the model’s underlying log-likelihoods. We apply a similar insight to a different problem: forecasting. We operationalize the log probability of gold-standard solutions not just as an explanatory tool, but as a direct covariate in a predictive scaling law for generative tasks, and we empirically demonstrate its superior stability for long-range extrapolation.

### v2

@@ 4.2 Predicting Pass Rates from Model Parameters and Pretraining Tokens
Figure 6: Scaling of GSM8K Pass Rates with Gold Reference Likelihood (Full Fit). Each panel fits Eqn. 6, $\log(\mathrm{pass}_{\mathcal{B}}@k)=\xi_{0}+K_{0}(k)\cdot[-\log\big(\mathrm{GoldProb}_{\mathcal{B}}\big)]^{\kappa(k)}$. Regressing the gold reference log-likelihoods against the pass rates produces a visually tighter fit with less residual scatter than the compute scaling law (cf. Fig. 1). As $k$ increases, the irreducible error $\xi_{0}(k)$ falls toward zero and the exponent $\kappa(k)$ increases, making the relationship a steeper, purer power law.
@@ 5 Fitting and Predicting Pass Rates from Gold Reference Likelihoods
Generative evaluations oftentimes contain not just problems and correct answers, but also “gold references”, high quality responses that reach the correct answer for the given problem. We calculated the average log-likelihood of these gold reference sequences to use to predict pass rates:
$\mathrm{GoldProb_{\mathcal{B}}}\quad\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}\quad\frac{1}{|\mathcal{B}|}\,\sum_{i\in\mathcal{B}}\,p_{\theta}(\text{Gold Reference}_{i}|\text{Problem}_{i}).$ (5)
For our gold reference likelihood scaling laws we assessed how well the benchmark $\mathrm{pass_{\mathcal{B}}@k}$ can be (1) fit and (2) predicted as a function of the log likelihoods of the gold references.
@@ 5.1 Fitting Pass Rates from Gold Reference Likelihoods
Figure 7: Predicting GSM8K Pass Rates from Gold Reference Likelihoods (Backtesting). Top Left: While the overall predictive error is comparable to the other laws, it is slightly lower for small $k$ and slightly higher for large $k$. Other Three Panels: The most striking result is the stability of the scaling law’s parameters. In stark contrast to the compute and parameters+tokens laws, whose parameters stabilized only when fit on models within $\mathord{\sim}2$ orders of magnitude of the target, the backtested estimates for $\widehat{\xi}_{0}(k)$, $\widehat{K}_{0}(k)$, and $\widehat{\kappa}(k)$ converge to their final values using models up to $\mathord{\sim}5$ orders of magnitude cheaper than the target. This remarkable stability suggests that the relationship between gold reference likelihoods and pass rates provides a robust signal for long-range forecasts.
Motivated by Schaeffer et al. (2024b), we tested whether this average likelihood of the gold responses can be accurately regressed to fit and predict benchmark pass rates $\mathrm{pass_{\mathcal{B}}}@k$.
@@
We find that this straightforwardly computable quantity serves as an exceptionally accurate predictor of benchmark pass rates. As shown in Figure 6, regressing against the negative log likelihood of the gold references produces a visually tighter fit than the pretraining compute law (Fig. 1). The residual scatter is reduced, and the relationship between $\log(\mathrm{pass_{\mathcal{B}}}@k)$ and $\log(\mathrm{GoldProb_{\mathcal{B}}})$ has a smaller reducible error and more closely resembles a power law. We next examined how the number of attempts per problem $k$ influences the scaling law’s parameters (Fig. 13). Consistent with our findings for the prior two scaling laws, the irreducible error term $\xi_{0}(k)$ collapses toward zero as $k$ increases and the exponent $\kappa(k)$ increases smoothly.
@@ 5.2 Predicting Pass Rates from Gold Reference Likelihoods
Using our backtesting approach once more, we evaluated how well the gold references scaling law can predict the performance of our most expensive model. For small $k$, the gold reference likelihoods scaling law has slightly lower relative error at predicting the most expensive model’s performance, but worsens for larger $k$ (Fig. 7 Upper Left). Remarkably, the gold reference law exhibits exceptional parameter stability. Unlike compute or parameter + token scaling laws, the backtested estimates for the gold reference scaling parameters stabilize when fitting on models up to $\mathord{\sim}5$ of magnitude cheaper than the target (Fig. 7, Other Three Panels) and for all values of $k$, suggesting this relationship is a robust signal for long-range forecasting.
6 How the Compute Law Emerges from the Parameters + Tokens Law, With Implications For Overtraining
The approximately equivalent performance between the compute scaling law (Eqn. 3) and the parameters + tokens scaling law (Eqn. 4) led us to ask whether a deeper connection might exist. We discovered that the compute scaling law is the compute-optimal envelope of the parameters+tokens scaling law. We summarize the insight here; see Appendix E for a deeper discussion.
Fix benchmark $\mathcal{B}$ and attempts per problem $k$. Consider a compute budget $C\approx c\,ND$. The best achievable error is obtained by minimizing the right-hand side of Eqn. 4 over all $(N,D)$ with $ND=C/c$. Evaluating at the minimizer yields Eqn. 3. The mapping of exponents and constants is:
$\alpha(k)=\Bigl(\tfrac{1}{\beta(k)}+\tfrac{1}{\gamma(k)}\Bigr)^{-1},\qquad E_{0}(k)=\mathcal{E}_{0}(k).$
Compute-Optimal Allocation: Along the envelope (i.e., when Eqn. 3 is tight), the optimal split of compute between parameters and tokens is
@@ 6 How the Compute Law Emerges from the Parameters + Tokens Law, With Implications For Overtraining
Thus the compute-optimal tokens and parameter shift with scale according to $\beta(k),\gamma(k)$; when $\beta(k)=\gamma(k)$, both grow like $C^{1/2}$.
Departures from the Envelope: Deviating from the compute-optimal allocation $(N^{*},D^{*})$ introduces a multiplicative penalty $\Phi(r)\geq 1$, where $r(C)=\tfrac{D^{*}(C)}{D}=\tfrac{N}{N^{*}(C)}$ is the misallocation ratio:
$\log\big(\operatorname{pass}_{\mathcal{B}}\text{@}k\big)(C,k)=E_{0}(k)+C_{0}(k)\cdot C^{-\alpha(k)}\cdot\underbrace{\Bigl[\tfrac{\gamma}{\beta+\gamma}\,r^{-\beta}+\tfrac{\beta}{\beta+\gamma}\,r^{\gamma}\Bigr]}_{\displaystyle\Phi(r;\beta(k),\gamma(k))\ \geq 1},$
Figure 8: Quantifying the Cost of Non-Optimal Scaling. We plot the “effective” compute multiplier (derived in Eq. LABEL:eqn:effective_compute_def) against the misallocation ratio $r(C)$. The peak at $r=1$ corresponds to the compute-optimal allocation; deviating from this optimum—either by overtraining ($r<1$, left) or undertraining ($r>1$, right) rapidly reduces the effective compute by orders of magnitude. Benchmark: GSM8K.
with equality iff $r\equiv 1$. Intuitively, $\Phi(r)$ represents a “tax” on performance caused by inefficient compute allocation: small deviations are benign ($\Phi=1+\tfrac{\beta\gamma}{2}\,(\log r)^{2}+O(\log r)^{3})$ ), but persistent off-ridge scaling significantly reduces “effective” pretraining compute:
$\frac{\text{Effective Pretraining Compute}}{\text{Pretraining Compute}}=\frac{C(r=1)}{C(r\neq 1)}$ (7)
$=\Phi(r;\beta(k),\gamma(k))^{-1/\alpha(k)}$
A useful case: if one scales $N$ and $D$ at a fixed ratio (i.e., without re-optimizing with $C$), the effective compute slope degrades to $\alpha_{\text{path}}=\frac{\min\{\beta,\gamma\}}{2}$.
Takeaways: (i) The compute law is not an independent phenomenology; it is the optimal-allocation shadow of the parameters + tokens law. (ii) Use the above allocation to stay on-envelope to maximize compute efficiency. (iii) Implementation constants relating $C$ to $ND$ shift $C_{0}(k)$, but do not change $\alpha(k)$. All statements hold pointwise in $k$.
Relationship to Overtraining: Our misallocation ratio $r(C)$ is the inverse of the overtraining multiplier $M$ studied by Gadre et al. (2024). Our results extend their overtraining scaling law to generative evaluations and the downstream metric $\mathrm{pass}_{\mathcal{B}}@k$ for $k\geq 1$.
7 How Scaling Laws Depend on the Benchmark: MATH vs GSM8K
We repeated the same scaling analyses for MATH (Hendrycks et al., 2021) as for GSM8K. Due to space constraints, we defer the detailed results to Appendix F and focus here on the salient similarities and differences between the benchmarks.
First, irreducible error reflects benchmark hardness. For GSM8K, the irreducible error $E_{0}(k)$ vanishes quickly, approaching zero by $k\approx 100$. In contrast, MATH retains significant irreducible error even at high sample counts: $E_{0}(k)\approx 0.45$ at $k=10^{4}$ (Figure 14). This quantifies the intuition that MATH is a significantly harder benchmark.
Second, harder benchmarks yield steeper power laws at high sample counts. While both benchmarks show that increasing $k$ steepens the scaling curve, MATH exhibits a more dramatic shift. The compute exponent $\alpha(k)$ for GSM8K rises to $\approx 0.38$ at $k=10^{4}$, whereas for MATH it rises to $\approx 0.58$. Because GSM8K performance saturates near 100% coverage, the returns on compute diminish. Conversely, because MATH performance remains far from perfect, the “headroom” allows for steeper scaling returns as compute increases.
Third, scaling exponents for MATH are sometimes more stable and sometimes less stable when backtesting. The compute-only scaling law’s $\alpha(k)$ and the parameters+tokens’s $\beta(k)$ stabilize across 3 orders of magnitude of compute, but the gold reference likelihood’s $\kappa(k)$ stabilizes across 2 order of magnitude and the parameters+tokens’s $\gamma(k)$ stabilizes across 1 order of magnitude. We do not know what exactly determines this - the hardness of the task, the distribution shift between training data and the task, or some other factor.
@@ 8 Discussion
Key Findings: We investigated pretraining scaling laws for generative evaluations of language models. Our work was motivated by OpenAI et al. (2024a)’s success, as well as by the insights of Hu et al. (2024). Our primary contribution is the proposal and rigorous backtesting of three distinct scaling laws for pass-at-$k$, utilizing (1) pretraining compute, (2) model parameters and pretraining tokens, and (3) gold reference log likelihoods as covariates. We demonstrate that the number of attempts per problem, $k$, acts as a critical hyperparameter in generative evaluation, offering a novel lever to control both the parameters of the scaling laws and their predictive reliability. Empirically, while all three laws exhibit comparable predictive accuracy, the gold reference likelihood law is uniquely stable, with parameters converging across nearly five orders of magnitude of compute. Finally, we establish a theoretical connection between the compute law and the parameters-and-tokens law, proving that the former emerges as the compute-optimal envelope of the latter. This derivation yields a dimensionless misallocation penalty, quantifying the efficiency loss when training deviates from the optimal parameter-token ratio.
Limitations: The primary limitation of this work is its empirical focus on a single model family (Pythia). As detailed in Section 2, this scope was a necessary constraint to enable a rigorous, high-resolution analysis (fitting 5-parameter laws requires dense checkpoints) while managing the substantial computational costs of generative sampling. Additionally, we focus exclusively on pass-at-$k$ metrics; other generative evaluation metrics and tasks warrant similar investigation.
Related Work: We defer a detailed discussion of related work to Appendix A.
Future Directions: Several avenues for future research are clear. First, subsequent work should determine whether predictive performance can be further improved using controlled pretrained models or more extensive sampling budgets. Second, OpenAI et al. (2024a) suggest that subsetting and/or stratifying problems enables better predictions; future work should isolate exactly which techniques yield the highest signal-to-noise. Third, the strong predictive power of gold reference log-likelihoods warrants theoretical investigation: given that next-token sampling is a branching process with likely exponentially many valid solution paths, it is not immediately obvious why the likelihood of the specific benchmark-provided gold reference correlates so strongly with the pass rate. Two key questions arise: (i) how generalizable is this relationship across diverse benchmarks? and (ii) to what extent does this signal remain robust under heavy optimization pressure? Fourth, while we held decoding strategies constant to isolate the effect of $k$, understanding how different sampling algorithms (e.g., top-$p$) and temperature interact with scaling laws remains an open and valuable question.
Acknowledgments
RS acknowledges support from Stanford Data Science and from the OpenAI Superalignment Fast Grant. NL is supported by the EPFL AI4science/AI Center program. BM acknowledge support by Schmidt Sciences, Stanford EDGE Scholar Fellowship. SK acknowledges support by NSF 2046795 and 2205329, the MacArthur Foundation, Stanford HAI, OpenAI and Google Inc.
@@ Appendix A Related Work
The study of neural scaling laws, which characterize how model performance predictably improves with resources, has become a cornerstone of modern large-scale machine learning (Hestness et al., 2017; Kaplan et al., 2020a; Hoffmann et al., 2022a; Roberts et al., 2022; Bahri et al., 2024). Foundational work demonstrated that pretraining loss follows a power law with respect to model size, dataset size, and training compute (Kaplan et al., 2020a; Henighan et al., 2020; Hoffmann et al., 2022a). These principles now guide the development of nearly all state-of-the-art language models (OpenAI et al., 2024a; Team et al., 2024; Grattafiori et al., 2024; DeepSeek-AI et al., 2025b; Qwen et al., 2025; Zhipu-AI et al., 2025). This research area has since expanded to explore scaling properties in diverse domains, including computer vision (Zhai et al., 2022), machine translation (Ghorbani et al., 2021), and even for optimizing inference-time compute rather than training (Wu et al., 2024; Snell et al., 2024). Despite this breadth, scaling laws for complex, multi-step generative tasks—which often represent the ultimate goal of language modeling—have remained comparatively underexplored due to the challenges of evaluation. Our work builds most directly on recent efforts to forecast performance on such generative evaluations. The GPT-4 Technical Report pioneered this direction by showing that the negative log pass rate on a subset of the HumanEval coding benchmark scaled predictably as a power law of pretraining compute (OpenAI et al., 2024a); they successfully predicted GPT-4’s performance by extrapolating from smaller models over approximately three orders of magnitude (OpenAI et al., 2024a). We adopt their proposed functional form for our compute-based law and extend their work by systematically analyzing how the number of attempts, $k$, reshapes the law’s parameters. Furthermore, our rigorous backtesting reveals the brittleness of this compute-only law, finding it reliable only across 1-2 orders of magnitude in our setting. Separately, our work shares a philosophical motivation with Hu et al. (2024), who investigate “infinite resolution” sampling for emergent abilities.
Relationship Between Pass-at-$k$ and Accuracy: In the context of language model evaluations, accuracy and pass@$k$ are fundamentally different. “Accuracy” is a likelihood comparison: does the correct multiple-choice answer have the highest probability relative to the incorrect alternative answers? (Gao et al., 2024; Schaeffer et al., 2024a). In contrast, pass@k measures the success of free-form generative sampling. To make the distinction clear, here are three obvious differences:
  1. 1.
    Accuracy is either binary 0 or 1, whereas pass@$k$ is a continuous quantity in $[0,1]$ and practically almost always in $(0,1)$. The two metrics have different output spaces.
  2. 2.
    As functions, the two metrics have different inputs: accuracy is a function of the correct answer and the incorrect answer, whereas pass@k does not depend on the incorrect answer.
  3. 3.
    Accuracy is computed using the exact sequence of the correct answer via teacher forcing. In comparison, pass@$k$ can sample any sequence of tokens.
Detailed Comparison with Gadre et al. (2024): On the empirical side, the 46 downstream tasks in Gadre et al. (2024) fall into two categories (see their Table 5): (1) pretraining losses (i.e., next-token prediction), (2) discriminative evaluations: measuring accuracy on multiple-choice tasks (e.g., MMLU). Our work focuses on a third, distinct category: generative evaluation, scored with pass@$k$.
On the theoretical side, our starting point is the standard separable ($(N,D)$) scaling law of Hoffmann et al. (2022a) ($L(N,D)\approx E+AN^{-\alpha}+BD^{-\beta}$). Gadre et al. (2024) show in their Appendix B that this form implies a compute power law with exponent ($\alpha\beta/(\alpha+\beta)$), and that departures from the compute‑optimal allocation modify only the prefactor as a function of an over‑training parameter (m). Our derivation in Appendix E is a direct generalization of this result to the relevant quantity for inference scaling ($\log\mathrm{pass}@k$) and to arbitrary ($k$): we retain the same ($(N,D)$) ansatz, derive the compute‑optimal envelope and its exponent ($\alpha(k)=(1/\beta(k)+1/\gamma(k))^{-1}$), and explicitly express how sub‑optimal choices of ($(N,D)$) enter as a multiplicative penalty depending on the ratio ($N/N^{*}$).